A bulb lights brightly drawing a current of 2A. how much charge runs through the bulb in 2 minutes?
1
Expert's answer
2011-02-07T06:48:04-0500
From to the definition
I = dQ/dt,& for the constant current I = Q/t.
we can find Q:
& Q = It = 2[A]*2*60 [sec] = 240 C